OpenStudy (anonymous):

First one to answer gets a medal. If Logx (1 / 8) = - 3 / 2, then x is equal to

OpenStudy (cwrw238):

1/8 = x ^ (-3/2) 1/8 = 1 / x^(3/2) x^(3/2) = 8 x = 4

OpenStudy (shadowfiend):

Sorry to hijack the question, but—just so we're clear, deleting insults doesn't mean you didn't say them. Please respect other members of OpenStudy, or we'll have to issue a suspension.